Find the cost per pound of a trail mix made from 10 lb of raisins that costs $4.20 per pound and 90 lb of granola that costs $5.60 per pound.
A. $4.32
B. $5.46
C. $6.07
D. $6.83

Answer by KnightOwlTutor(293)

The amounts are equal so that you have a total of 100 lbs on each side of the equation
On the one side of the equation we are finding the total cost of the individual componets of the trail mix. We found the individual cost of the raisins and granola used to make the trail mix.
10(4.20)+90(5.60)=(10+90)X
On the other side of the equation we know that we made 100 lbs of the trail mix but we don't know the cost that is our unknown X.
Use the distributive property
42+504=100X
546=100X
X=5.46
The answer is B. $5.46/lb